OpenStudy (anonymous):

Please help!!! WILL FAN + MEDAL Can you describe multiplying and dividing signed numbers in the real world??

OpenStudy (sloppycanada):

Ordering pizza. So one large pizza costs 5.00 and since I want three of them, 5 x 3 = 15. I'll be paying 15 dollars for 3 large pizzas

OpenStudy (anonymous):

What about the signed numbers part?

OpenStudy (anonymous):

@imqwerty

OpenStudy (jadedry):

I assume that signed numbers means you're dealing with numbers that have a plus or minus sign attached to them?

OpenStudy (anonymous):

Yes

OpenStudy (jadedry):

So I assume this example would do: You have $100 in your bank account. You and your friend want to purchase something that cost's $120. You draw $120, and divide the debt with your friend. You both have -$10 / Owe the bank $10

imqwerty (imqwerty):

ok remember these rules-> +ve ---> a positive number -ve--> a negative number multiplication rules- \[(+ve) \times (+ve)=+ve\]\[(+ve) \times (-ve)=-ve\]\[(-ve) \times (+ve)=-ve\]\[(-ve) \times (-ve)=+ve\] division rules-\[\frac{ +ve }{ +ve }=+ve\]\[\frac{ +ve }{ -ve }=-ve\]\[\frac{ -ve }{ +ve }=-ve\]\[\frac{ -ve }{ -ve }=+ve\] or u can just remember this- whenever u have 2positives or 2negatives in a multiplication or division then u always get a positive number but if u have a positive and a negative then u will always get a negative number

OpenStudy (anonymous):

Ahh thank you!!! x

OpenStudy (jadedry):

@imqwerty That's definitely more thorough than my answer haha. Good one! c:

OpenStudy (anonymous):

@Jadedry your explanation of a real world situation was very helpful! thank you both. x

imqwerty (imqwerty):

np :)